# Stability of Non-Hyperbolic Sets

### v1

@@ 8 Finite Collet-Eckmann Condition
    $\frac{1}{4K^{6}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n+n_{0}}\left(\lambda^{\theta}\mu\right)^{n}\geq\rho^{N+k},\text{ }\forall 1\leq k\leq M$
@@
    $\frac{1}{4K^{6}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{0}}\left(\frac{\lambda^{\theta}\mu^{2}}{\rho}\right)^{n}\geq\rho^{N+M}.$
    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P6 proves
    $\left|Df^{T+n+N+n_{0}+M}_{x}v\right|\geq\frac{C}{4K^{6}}\cos\varphi\mu^{n_{0}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|.$ (8.13)
@@
    and by (8.12),
    $\text{d}_{\text{v}}\left(f^{T+n+N+n_{0}}(x),q_{2}\right)<<1.$
    The previous two inequalities together with $\text{IH}_{H}$ gives
    $\angle\left(Df^{T+n+N+n_{0}+M}_{x}v,e_{1}\right)>\frac{\beta}{2}.$
@@
    $v^{(6)}=Df^{T+n+N+n_{0}+M}_{x}v=\left(\begin{matrix}v^{(6)}_{1}\\ v^{(6)}_{2}\end{matrix}\right)$,
@@
    $\frac{v^{(6)}_{2}}{\left|v^{(6)}\right|}\geq\sin\frac{\beta}{2}.$ (8.14)
@@
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+nN+n_{0}+M+1$ up to time $T+n+N+n_{0}+M+\overline{\ell}$. Indeed, by $\text{IH}_{Q}$, (8.14) and (8.13), for all $1\leq k\leq\overline{\ell}$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P7:
      $\left|Df^{T+n+N+n_{0}+M+k}_{x}v\right|$ $\geq$ $\mu^{k}\sin\frac{\beta}{2}\left|Df^{T+n+N+n_{0}+M}_{x}v\right|$
      $\geq$ $\frac{C}{4K^{6}}\cos\varphi\sin\frac{\beta}{2}\mu^{n_{0}+k}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|$
      $\geq$ $C\rho^{T+n+N+n_{0}+M+k}\left|v\right|.$
@@
    $\frac{1}{4K^{6}}\cos\varphi\sin\frac{\beta}{2}\left(\frac{\mu}{\rho}\right)^{n_{0}+k}\left(\frac{\lambda^{\theta}\mu^{2}}{\rho}\right)^{n}\geq\rho^{N+M}.$
    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. Denote by
    $v^{(7)}=Df^{T+n+N+n_{0}+M+\overline{\ell}}_{x}v=\left(\begin{matrix}v^{(7)}_{1}\\ v^{(7)}_{2}\end{matrix}\right)$,
    then, by $\text{IH}_{Q}$ and by (8.14), we get
    $\frac{v^{(7)}_{1}}{v^{(7)}_{2}}=\left(\frac{\lambda}{\mu}\right)^{\overline{\ell}}\frac{v^{(6)}_{1}}{v^{(6)}_{2}}\leq\left(\frac{\lambda}{\mu}\right)^{\overline{\ell}}\cot\frac{\beta}{2}<\tan\varphi,$
    with the last inequality achieved by taking $\overline{\ell}$ large enough.
In conclusion, the inequality above proves $\text{IH}_{\tilde{C}}-$(ii) and $\text{IH}_{\tilde{C}}-$(i) is achieved by putting together $\text{IH}_{\tilde{C}}-$(i)P1, $\text{IH}_{\tilde{C}}-$(i)P2, $\text{IH}_{\tilde{C}}-$(i)P3, $\text{IH}_{\tilde{C}}-$(i)P4, $\text{IH}_{\tilde{C}}-$(i)P5, $\text{IH}_{\tilde{C}}-$(i)P6 and $\text{IH}_{\tilde{C}}-$(i)P7.
9 The main objects: the lamination and the invariant Cantor set
@@ 9.1 The lamination
Consider the $1$-tangency strip $S$ given by $S=[-t_{0},t_{0}]\times[-a_{0},a_{0}]$ with tangency line $b=\left\{a=0\right\}$. Take $U=S$. In Subsection 2.1, we defined a family $F$ which is an unfolding of the tangency $b$ and each map $f\in S$ has a $(1,N_{1},M_{1})$ dynamical partition with $N_{1}=N$ and $M_{1}=M$, where $N$ and $M$ are defined in Definition 2, property $(f10)$.
@@
$\mathfrak{S}_{2}=\left\{S_{j,2}\right\}_{j=1}^{\infty}$
where each $S_{j,2}$ is a $n_{2}+j$-tangency strip contained in $S$ whose existence is guaranteed by Proposition 7. By applying Proposition 7 inductively, for all $g\geq 3$ one can define a sequence of tangency strips
$\mathfrak{S}_{g}=\left\{S_{j,g}\right\}_{j=1}^{\infty}$
@@
  • $\forall S_{j,g}\in\mathfrak{S}_{g}$ there exists a unique $S_{i,g-1}\in\mathfrak{S}_{g-1}$ with $S_{j,g}\subset S_{i,g-1}$.
We are now ready to define the lamination
$\mathfrak{L}=\bigcap_{g=1}^{\infty}\bigcup_{j=1}^{\infty}S_{j,g}.$ (9.1)
A leave of the lamination is defined as
$\ell=\bigcap_{g=1}^{\infty}S_{g}$ (9.2)
where, for all $g$, $S_{g}\in\mathfrak{S}_{g}$ and $S_{g}\subset S_{g-1}$. Observe that, by Equation (7.1), $\mathfrak{L}$ is a ${\mathcal{C}}^{1}$} lamination, i.e. each leave is the graph of a ${\mathcal{C}}^{1}$} function.
The following proposition proves Theorem A-(v).
Proposition 9.
Every leaves of $\mathfrak{L}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$. Morever if $t\in[-t_{0},t_{0}]$, the closure of the set $L_{t}=\left\{(t,a)\in\mathfrak{L}\right\}$ is a Cantor set.
Proof.
The proof follows from the construction of the lamination, and Proposition 7 and Proposition 4. The main point is that each strip contains leaves corresponding to maps with a new homoclinic tangency and these leaves are graphs of global curves. ∎
@@ 9.2 The invariant Cantor set
The following discussion is consequence of Proposition 7.
Given a leave of the lamination $\ell$ and a map $f\in\ell$, one can associate a sequence $\left\{\left(b_{g},U_{g},\mathfrak{D}_{g}\right)\right\}_{g=1}^{\infty}$ and a sequence of numbers $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ such that, for all $g$, $(b_{g},U_{g})$ is the $(1,N_{g},M_{g})$ descendant of $S_{g-1}$ and $\mathfrak{D}_{g}$ is the $(1,N_{g},M_{g})$ dynamical partition of $f$. The sequence $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ is called the type of $\ell$.
@@
$A(f)=\bigcap_{g=1}^{\infty}\mathcal{D}_{g}=\bigcap_{g=1}^{\infty}\bigcup_{D_{g}\in\mathfrak{D}_{g}}{D}_{g}$
where $\mathcal{D}_{g}$ is the union of the elements of $\mathfrak{D}_{g}$. This set is going to play a fundamental role in our research. Let us begin by studying its properties. Following Proposition 7, for all $g$, let $X_{g}$ be the $(1,N_{g},M_{g})$ directed graph and let $w_{g}:X_{g+1}\to X_{g}$ be the projection and $w_{g^{\prime}g}:X_{g}\to X_{g^{\prime}}$ be
@@ 9.3 Properties of the set AA
In this subsection we collect properties of the set $A$. In particular, the following lemma proves Theorem A-(i).
Lemma 22.
Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. Then the set $A(f)$ is a Cantor set.
@@
The set $A(f)$ is compact because it is the intersection of a nested sequence of compact sets and it is totally disconnected because the $\text{mesh}(\mathfrak{D}_{g})$ goes to zero when $g$ goes to infinity.
It is left to prove that $A(f)$ is perfect. Let us start by observing that, by Remark 14, each set $D_{g}\in\mathfrak{D}_{g}$ contains at least two sets $D_{g+1},D^{\prime}_{g+1}\in\mathfrak{D}_{g+1}$.
@@
In the following lemma we describe the possible limit behavior of points in $A$. We denote by $\omega(x)$, the $\omega$-limit set of the point $x$. In particular, it implies that the set $A$ is non-hyperbolic.
Recall that $q_{2}$ is the point with a transversal homoclinic intersection. Let
@@
The proof of the following lemmas relies on the directed graphs introduced before. We will need the following notation. Denote the corresponding prime loops in $X_{g}$ respectively by $\lambda^{q}_{g}$, $\lambda^{h}_{g}$, and $\lambda^{c}_{g}$. We will use frequently the following observation. If a loop in $X_{g}$ consists purely of $h$- and $q$-loops of $X_{g}$ then $w_{g^{\prime}g}(\lambda)$ consists purely of $h$- and $q$-loops of $X_{g^{\prime}}$.
Lemma 23.
Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. Each $x\in A(f)$ satisfies one of the following:
@@
$w_{g-1}\left(\overrightarrow{x_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$
@@
Hence, there exists $t$ such that $f^{t}\left(D_{g_{0}+2}\right)\subset\tilde{D}_{g_{0}}$ and in particular $f^{t}(x)\in\tilde{D}_{g_{0}}\subset U$.
@@
The proof relies on the following claim. Let $g_{0}\leq g^{\prime}<g$. There exists $t_{g^{\prime}}\geq 0$ such that for every $g>g^{\prime}$
$w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})=\lambda_{g^{\prime}}^{q}\cup\lambda_{g^{\prime}}^{h},$
where $y=f^{t_{g^{\prime}}}(x)$. In other words, the loop $w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})$ is a concatenation of multiple times the $q$-loop and $h$-loop of $X_{g^{\prime}}$.
The proof of this Claim is by induction. First observe
$w_{g^{\prime}}(\overrightarrow{x_{g^{\prime}+1}q_{g^{\prime}+1}})=\gamma\lambda,$
@@
where $\lambda$ is a loop which is a concatenation of multiple $\lambda_{g}^{q}$ and one $\lambda_{g}^{h}$. This is true because $w_{g}(\overrightarrow{y_{g+1}q_{g+1}})$ cannot contain the $c$-loop of $X_{g}$, we are in the second case. The induction assumption implies that
@@
$\text{Orbit}(f^{t_{g}^{\prime}}(x))\subset Q_{g^{\prime}}\cup\bigcup_{j=1}^{M_{g^{\prime}}-1}H_{g^{\prime}}(j).$
@@
$\omega(x)\subset\bigcap_{g\geq g_{0}}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right)$
where for all $g$ and for all $j$, $Q_{g}$ and $H_{g}(j)$ are the sets in $\mathfrak{D}_{g}$ as described in Definition 7. In order to complete the proof one should establish that
$\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right)=O_{2}.$
One inclusion is trivial. Let us concentrate on the other one and for this purpose let us take a point
$x\in\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right).$
@@
$x\in\bigcap_{g\geq\tilde{g}}\bigcup_{j=1}^{M_{g}-1}H_{g}(j).$
@@
$x\in H_{g}(j_{x}).$
In particular, there exists a number $t_{g}$ such that
$f^{t_{g}}(q_{2})\in H_{g}(j_{x}).$
Because, $f^{t}(q_{2})\to p$ when $|t|\to\infty$ we have that $t_{g}$ is bounded. Observe also that $f^{t_{g}}(q_{2})$ is close to $x$ because the diameter of $H_{g}(j_{x})$ goes to zero when $g$ goes to infinity. It follows that $f^{t}(q_{2})=x$ form some $t\in\mathbb{Z}$. Hence, $x\in O_{2}$ and the proof is complete. ∎
A natural question at this point is: does a point with dense orbit exist? The answer is yes and the proof is the following lemma which also proves Theorem A-(ii).
Lemma 24.
Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. There exists a point $c\in A(f)$ such that $\omega(c)=A(f).$
@@
Fix $C_{1}\in\mathfrak{D}_{1}$ and consider $c_{1}=\pi_{1}(C_{1})\in X_{1}$ where $\pi_{1}$ is the projection of $\mathcal{D}_{1}$ on $X_{1}$. We want to prove that for infinitely many $g$ there is an element $c_{g}$ in the $c$-loop of $X_{g}$ such that
@@
Suppose we have constructed $c_{1},\dots,c_{g-1}$. Consider the unique path $\gamma=\overrightarrow{c_{g-1}q_{g-1}}$ contained in the $c$-loop of $X_{g-1}$. Take the first vertex $c_{g}$ in the $c$-loop of $X_{g}$ with the property that $w_{g-1}(c_{g})=c_{g-1}$. Observe that
$w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)=\gamma\circ\lambda$
where $\lambda$ is a loop and, by Remark 14 $\lambda$ contains all primary loops of $X_{g-1}$. In particular $w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$.
For all $g$, consider the set of $\mathfrak{D}_{g}$ defined by $\pi_{g}(C_{g})=c_{g}$, where $\pi_{g}$ is the projection of $D_{g}$ on $X_{g}$. Our selected point is
@@
Following the proof of Lemma 23 one gets that $\omega(c)=A(f)$. ∎
The following lemma proves Theorem A-(iii).
Lemma 25.
Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. There exists a point $b\in A(f)$ such that $\omega(b)=O_{2}.$
@@
$w_{g-1}(\overrightarrow{x_{g}q_{g}})=\overrightarrow{x_{g-1}q_{g-1}}\lambda,$
@@
$w_{g_{0}g}(\overrightarrow{y_{g+1}q_{g+1}})=w_{g_{0}g}(\overrightarrow{y_{g}q_{g}})w_{g_{0}g}(\lambda).$
@@
$\omega(x)\subset Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)$

### v3

@@ 8 Finite Collet-Eckmann Condition
    $\frac{1}{4K^{6}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}+n_{0,k}^{(j)}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}\geq\rho^{N+s},\text{ }\forall 1\leq s\leq M$
@@
    $\frac{1}{4K^{6}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{0,k}^{(j)}}\left(\frac{\lambda^{\theta}\mu^{2}}{\rho}\right)^{n_{k}^{(j)}}\geq\rho^{N+M}.$
    This follow from the condition imposed in (8)-(oo) and taking $n_{k}^{(j)}$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P6 proves
    $\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M}_{x}v\right|\geq\frac{C}{4K^{6}}\cos\varphi\cdot\mu^{n_{0,k}^{(j)}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n_{k}^{(j)}}\left|v\right|.$ (8.0.15)
@@
    and by (7.4.3),
    $\text{dist}\left(f^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}}(x),q_{2}\right)<<1.$
    The previous two inequalities, (8.0.14) together with $\text{IH}_{H}$ gives
    $\angle\left(Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M}_{x}v,e_{1}\right)>\frac{\beta}{2}.$
@@
    $v^{(6)}=Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M}_{x}v=\left(\begin{matrix}v^{(6)}_{1}\\ v^{(6)}_{2}\end{matrix}\right)$,
@@
    $\frac{v^{(6)}_{2}}{\left|v^{(6)}\right|}\geq\sin\frac{\beta}{2}.$ (8.0.16)
@@
    Now we prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M+1$ up to time $\tilde{T}=T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M+\underline{\ell}_{k}^{(j)}$. Indeed, by $\text{IH}_{Q}$, (8.0.16) and (8.0.15), for all $1\leq s\leq\underline{\ell}_{k}^{(j)}$, we get,
    $\text{IH}_{\tilde{C}}\text{-(i)P7: }\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M+s}_{x}v\right|$ $\geq$ $\mu^{s}\sin\frac{\beta}{2}\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M}_{x}v\right|$
    $\geq$ $\frac{C}{4K^{6}}\cos\varphi\cdot\sin\frac{\beta}{2}\mu^{n_{0,k}^{(j)}+s}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n_{k}^{(j)}}\left|v\right|$
    $\geq$ $C\rho^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M+s}\left|v\right|.$
@@
    $\frac{1}{4K^{6}}\cos\varphi\cdot\sin\frac{\beta}{2}\left(\frac{\mu}{\rho}\right)^{n_{0,k}^{(j)}+s}\left(\frac{\lambda^{\theta}\mu^{2}}{\rho}\right)^{n_{k}^{(j)}}\geq\rho^{N+M}.$
    This follow from the condition imposed in (8)-(oo) and taking $n_{k}^{(j)}$ large enough. Denote by
    $v^{(7)}=Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M+\underline{\ell}_{k}^{(j)}}_{x}v=\left(\begin{matrix}v^{(7)}_{1}\\ v^{(7)}_{2}\end{matrix}\right)$,
    then, by $\text{IH}_{Q}$ and by (8.0.16), we get
    $\frac{v^{(7)}_{1}}{v^{(7)}_{2}}=\left(\frac{\lambda}{\mu}\right)^{\underline{\ell}_{k}^{(j)}}\frac{v^{(6)}_{1}}{v^{(6)}_{2}}\leq\left(\frac{\lambda}{\mu}\right)^{\underline{\ell}_{k}^{(j)}}\frac{1}{\sin\beta/2}<\tan\varphi,$
    with the last inequality achieved by taking $j$ large enough which implies that $\underline{\ell}_{k}^{(j)}$} is large enough.
In conclusion, the inequality above proves $\text{IH}_{\tilde{C}}-$(ii) and $\text{IH}_{\tilde{C}}-$(i) is achieved by putting together $\text{IH}_{\tilde{C}}-$(i)P1, $\text{IH}_{\tilde{C}}-$(i)P2, $\text{IH}_{\tilde{C}}-$(i)P3, $\text{IH}_{\tilde{C}}-$(i)P4, $\text{IH}_{\tilde{C}}-$(i)P5, $\text{IH}_{\tilde{C}}-$(i)P6 and $\text{IH}_{\tilde{C}}-$(i)P7. ∎
9 The lamination and the invariant Cantor set
@@ 9.1 The lamination
Consider the $1$-tangency strip $S$ given by $S=[-t_{0},t_{0}]\times[-a_{0},a_{0}]$ with tangency line $b=\left\{a=0\right\}$. Take $U=S$. In Subsection 2.1, we defined a family $F$ which is an unfolding of the tangency $b$ and each map $f\in S$ has a $(1,N_{1},M_{1})$ dynamical partition with $N_{1}=N$ and $M_{1}=M$, where $N$ and $M$ are defined in Definition 2.1.1, property $(f8)$.
@@
$\mathfrak{S}_{2}=\left\{S_{j,2}\right\}_{j=1}^{\infty},$
where each $S_{j,2}$ is a $n_{2}+j$-tangency strip contained in $S$ whose existence is guaranteed by Proposition 7.4.1. By applying Proposition 7.4.1 inductively, for all generations $g\geq 3$ one can define a sequence of collections of tangency strips
$\mathfrak{S}_{g}=\left\{S_{j,g}\right\}_{j=1}^{\infty},$
@@
  • $\forall S_{j,g}\in\mathfrak{S}_{g}$ there exists a unique $S_{i,g-1}\in\mathfrak{S}_{g-1}$ with $S_{j,g}\subset S_{i,g-1}$,
  • $\forall j,$ $S_{j,g}$ is bounded by the graph of two functions $s_{j,g}^{+},s_{j,g}^{-}:[-t_{0},t_{0}]\to\mathbb{R}$ such that,
    $\left\lVert s_{j,g}^{+}-s_{j,g}^{-}\right\rVert_{0}\leq\frac{1}{2^{g}},$
    and if $S_{j,g}\subset S_{i,g-1}$, then
    $\left\lVert\frac{ds_{j,g}^{\pm}}{dt}-\frac{ds_{i,g-1}^{\pm}}{dt}\right\rVert_{0}\leq\frac{1}{2^{g}}.$
This last property relies on (7.4.2). Observe that $\cup\mathfrak{S}_{g}\subset\cup\mathfrak{S}_{g-1}$. We are now ready to define the lamination
$\mathfrak{L}=\bigcap_{g=1}^{\infty}\bigcup_{j=1}^{\infty}S_{j,g}\subset[-t_{0},t_{0}]\times[0,a_{0}].$ (9.1.1)
The connected components of $\mathfrak{L}$ are called leafs of the lamination. Namely, a leaf of the lamination has the form
$\ell=\bigcap_{g=1}^{\infty}S_{g},$ (9.1.2)
where, for all $g$, $S_{g}\in\mathfrak{S}_{g}$ and $S_{g}\subset S_{g-1}$. Observe that $\mathfrak{L}$ is a ${\mathcal{C}}^{1}$} lamination, i.e. each leaf is the graph of a ${\mathcal{C}}^{1}$} function, $\lim_{q\to\infty}s^{+}_{g}$ where the graph of $s^{+}_{g}$ is the upper boundary of $S_{g}$.
The following proposition proves Theorem A-(v). The proof is a consequence of the construction of the lamination and of the fact that the strips of the same generation are pairwise disjoint.
Proposition 9.1.3.
Every leaf of $\mathfrak{L}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$. Morever if $t\in[-t_{0},t_{0}]$, the closure of the set $L_{t}=\left\{(t,a)\in\mathfrak{L}\right\}$ is a Cantor set.
Remark 9.1.4.
As in the proof of Proposition 8.0.4 by taking all $n_{g}^{(j)}$ large enough we may also assume that, for all $g$, $N_{g+1},M_{g+1}\geq 32\max\{N_{g},M_{g}\}$ and $N_{g},M_{g}\geq 3$.
The following Remark proves Theorem A-(viii)
Remark 9.1.5.
Observe that, by construction, given an open neighborhood of $f\in\mathfrak{L}$, it contains in parameter space a curve $b$ of secondary tangencies. Hence, by Remark 6.2.6, such open neighborhood contains maps with infinitely many sinks.
@@ 9.2 The invariant Cantor set
Given a leaf $\ell=\cap S_{g}$ of the lamination $\mathfrak{L}$ and a map $f\in\ell$, one can associate a sequence $\left\{\left(b_{g},U_{g},\mathfrak{D}_{g}\right)\right\}_{g=1}^{\infty}$ and a sequence of numbers $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ such that, for all $g$, $(b_{g},U_{g})$ is the $(1,N_{g},M_{g})$ descendant of $S_{g-1}$ and $\mathfrak{D}_{g}$ is the $(1,N_{g},M_{g})$ dynamical partition of $f$. The sequence $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ is called the type of $\ell$.
@@
$A(f)=\bigcap_{g=1}^{\infty}\mathcal{D}_{g}=\bigcap_{g=1}^{\infty}\bigcup_{D\in\mathfrak{D}_{g}}{D}$
where $\mathcal{D}_{g}$ is the union of the elements of $\mathfrak{D}_{g}$. This set is going to play a fundamental role. Let us begin by studying its properties. Following Proposition 7.4.1, for all $g$, let $X_{g}$ be the $(1,N_{g},M_{g})$ directed graph, let $w_{g}:X_{g+1}\to X_{g}$ be the projection and $w_{g^{\prime}g}:X_{g}\to X_{g^{\prime}}$ be
@@ 9.3 Properties of the set AA
In this subsection we collect properties of the set $A$.
Lemma 9.3.1.
Let $\ell$ be a leaf of the lamination $\mathfrak{L}$ and let $f\in\ell$. Then the set $A(f)$ is a Cantor set.
@@
The set $A(f)$ is compact because it is the intersection of a nested sequence of compact sets and it is totally disconnected because the $\text{mesh}(\mathfrak{D}_{g})$ goes to zero when $g$ goes to infinity. It is left to prove that $A(f)$ is perfect. First observe that, by Remark 7.4.4, each set $D_{g}\in\mathfrak{D}_{g}$ contains at least two sets $D_{g+1},D^{\prime}_{g+1}\in\mathfrak{D}_{g+1}$, see Figure 12.
@@
In the following lemma we describe the possible limit behavior of points in $A$. In particular, it implies that the set $A$ is non-hyperbolic.
We denote by $\omega(x)$, the $\omega$-limit set of the point $x$. Recall that $q_{2}$ is the point with a transversal homoclinic intersection. Let
@@
The proof of the following lemmas relies on the directed graphs introduced before. We will need the following notation. Denote the corresponding prime loops in $X_{g}$ respectively by $\lambda^{q}_{g}$, $\lambda^{h}_{g}$, and $\lambda^{c}_{g}$. We will use frequently the following observation: if a loop in $X_{g}$ consists purely of $h$- and $q$-loops of $X_{g}$ then $w_{g^{\prime}g}(\lambda)$ consists purely of $h$- and $q$-loops of $X_{g^{\prime}}$.
Lemma 9.3.2.
Let $\ell$ be a leaf of the lamination $\mathfrak{L}$ and let $f\in\ell$. Each $x\in A(f)$ satisfies one of the following:
@@
$w_{g-1}\left(\overrightarrow{x_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$ (9.3.3)
@@
Hence, there exists $t$ such that $f^{t}\left(D_{g_{0}+2}\right)\subset\tilde{D}_{g_{0}}$ and in particular $f^{t}(x)\in\tilde{D}_{g_{0}}\subset U$. In particular we showed that under if (9.3.3) holds for infinitely many $g$ then $\omega(x)=A(f)$.
@@
The proof relies on the following claim. Let $g_{0}\leq g^{\prime}$. There exists $t_{g^{\prime}}\geq 0$ such that the point $y=f^{t_{g^{\prime}}}(x)$ has the property that for every $g>g^{\prime}$,
$w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})=\lambda_{g^{\prime}}^{q}\cup\lambda_{g^{\prime}}^{h}.$
In other words, the path $w_{g^{\prime}g}(\overrightarrow{y_{g}q_{g}})$ is a concatenation of multiple times the $q$-loop and $h$-loop of $X_{g^{\prime}}$. The proof of this Claim is by induction. First observe
$w_{g^{\prime}}(\overrightarrow{x_{g^{\prime}+1}q_{g^{\prime}+1}})=\lambda\gamma,$
@@
where $\lambda$ is a loop which is a concatenation of multiple $\lambda_{g}^{q}$ and one $\lambda_{g}^{h}$. This is true because $w_{g}(\overrightarrow{y_{g+1}q_{g+1}})\subset w_{g}(\overrightarrow{x_{g+1}q_{g+1}})$ which cannot contain the $c$-loop of $X_{g}$, we are in the second case. The induction assumption implies that
@@
$\text{Orbit}(f^{t_{g^{\prime}}}(x))\subset Q_{g^{\prime}}\cup\bigcup_{j=1}^{M_{g^{\prime}}-1}H_{g^{\prime}}^{j}.$
@@
$\omega(x)\subset\bigcap_{g\geq g_{0}}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}\right)$
where for all $g$ and for all $j$, $Q_{g}$ and $H_{g}^{j}$ are the sets in $\mathfrak{D}_{g}$ as described in Definition 7.1.1. In order to complete the proof one should establish that
$\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}\right)=O_{2}.$
One inclusion is trivial, since each $h$-loop contains $q_{2}$. Let us concentrate on the other one and for this purpose let us take a point
$x\in\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}\right).$
@@
$x\in\bigcap_{g\geq\tilde{g}}\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}.$
@@
$x\in H_{g}^{j_{x}}.$
In particular, there exists a number $t_{g}\in\mathbb{Z}$ such that
$f^{t_{g}}(q_{2})\in H_{g}^{j_{x}}.$
Because, $f^{t}(q_{2})\to p$ when $|t|\to\infty$ we have that $t_{g}$ is bounded. Observe also that $f^{t_{g}}(q_{2})$ is close to $x$ because the diameter of $H_{g}^{j_{x}}$ goes to zero when $g$ goes to infinity. It follows that $f^{t}(q_{2})=x$ for some $t\in\mathbb{Z}$. Hence, $x\in O_{2}$ and the proof is complete. ∎
A natural question at this point is: does a point with dense orbit exist? The answer is yes. The following lemma also confirms Theorem A-(ii).
Lemma 9.3.4.
Let $\ell$ be a leaf of the lamination $\mathfrak{L}$ and let $f\in\ell$. There exists a point $c\in A(f)$ such that $\omega(c)=A(f).$
@@
Chose $C_{1}\in\mathfrak{D}_{1}$ such that $c_{1}=\pi_{1}(C_{1})\in\lambda^{c}_{1}$ where $\pi_{1}$ is the projection of $\mathcal{D}_{1}$ on $X_{1}$. We want to prove that for all $g$ there is an element $c_{g}$ in the $c$-loop of $X_{g}$ such that
@@
Suppose we have constructed $c_{1},\dots,c_{g-1}$. Take the first vertex $c_{g}$ in the $c$-loop of $X_{g}$ with the property that $w_{g-1}(c_{g})=c_{g-1}$. Observe that
$w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)=\lambda\overrightarrow{c_{g-1}q_{g-1}}$
where $\lambda$ is a loop and, by Remark 7.4.4 $\lambda$ contains all primary loops of $X_{g-1}$. In particular $w_{g-1}\left(\overrightarrow{c_{g}q_{g}}\right)\supset\lambda^{c}_{g-1}$.
For all $g$, consider the set of $C_{g}$ defined by $\pi_{g}(C_{g})=c_{g}$, where $\pi_{g}$ is the projection of $\mathcal{D}_{g}$ on $X_{g}$. Our selected point is
@@
Observe that $c$ satisfies (9.3.3) for all $g$. Hence, $\omega(c)=A(f)$. ∎
The following lemma confirms Theorem A-(iii).
Lemma 9.3.5.
Let $\ell$ be a leaf of the lamination $\mathfrak{L}$ and let $f\in\ell$. There exists a point $b\in A(f)$ such that $\omega(b)=O_{2}.$
@@
$w_{g-1}(\overrightarrow{x_{g}q_{g}})=\lambda\overrightarrow{x_{g-1}q_{g-1}},$
@@
$w_{g_{0}g}(\overrightarrow{y_{g+1}q_{g+1}})=w_{g_{0}g}(\lambda)w_{g_{0}g}(\overrightarrow{y_{g}q_{g}}).$
@@
$\omega(x)\subset Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}$